OpenStudy (anonymous):

help me what is the solution of x^2+6x=-5

OpenStudy (solomonzelman):

\(\Large \color{blue}{ \bf x^2+6x=-5}\) \(\Large \color{blue}{ \bf +5~~~~~~~~~~~+5}\) \(\Large \color{blue}{ \bf x^2+6x+5=0}\) factor.... \(\Large \color{blue}{ \bf (x+1)(x+5)=0}\) ca you take it from here?

OpenStudy (anonymous):

not really

OpenStudy (solomonzelman):

\(\Large \color{blue}{ \bf (x+1)(x+5)=0 }\) so either \(\Large \color{blue}{ \bf (x+1)=0 }\) or \(\Large \color{blue}{ \bf (x+5)=0 }\) solve each equation and tell me what you get for x.

OpenStudy (anonymous):

ok